# MEDSAGE: Enhancing Robustness of Medical Dialogue Summarization to ASR Errors with LLM-generated Synthetic Dialogues

### v1

@@ 1 Introduction
Automatic Speech Recognition (ASR) (Yu and Deng 2016) is the task of transcribing speech signals into text, enabling a wide range of applications from voice-activated assistants to automated customer service systems. ASR systems significantly aid various downstream tasks such as dialogue summarization (Zhong et al. 2022), where the goal is to distill key information from spoken interactions. However, errors introduced by ASR systems can degrade the performance of these summarization tasks (Li et al. 2014; Guo et al. 2024), which limits their application in high-stake domains where correctness of the summaries is important. The synthesis of Electronic Medical Records (EMRs) from doctor-patient dialogues (Krishna et al. 2021) is among such summarization tasks where accuracy is critical. Here, ASR errors can lead to inaccuracies in transcribing clinical terms, medication, or procedure names, resulting in erroneous medical notes (Hodgson and Coiera 2016), which can lead to misdiagnosis, incorrect treatment plans, and potentially harmful patient outcomes.
One way to improve dialogue summarization is to improve the ASR systems. However, this requires large amounts of supervised data (Radford et al. 2022), which is often unavailable in the healthcare domain due to privacy and ethical concerns surrounding the recording of doctor-patient conversations. Consequently, practical clinical summarization systems often treat ASR systems as black boxes, mandating that overall improvements must arise either from post-processing ASR outputs as a separate step, or making down-stream methods robust to ASR errors. Recent works propose post-processing ASR outputs using LLMs to correct erroneous ASR transcripts (Radhakrishnan et al. 2023; Bai et al. 2024). Nonetheless, based on prior studies, using prompting techniques to clear noise only proves effective only when using LLMs of large sizes, typically those that exceed 100B parameters (Yang et al. 2023). Moreover, LLMs are prone to hallucinations (Maynez et al. 2020), which can introduce irrelevant symptoms or medication names into the dialogue transcript, potentially degrading summarization quality while attempting to clear ASR noise.
@@
Our experimental evaluation reveals that large language models (LLMs) are effective noise modelers, capable of producing errors similar to those found in transcripts produced by ASR. This is evident from the qunatitative and qualitative similarities between the word error profiles of the synthetic dialogues we generate and actual ASR dialogue transcripts. Such similarity also reflects on the downstream summarization performance, mirroring the performance drop caused by ASR errors. Lastly, when we utilize these synthetic noisy dialogues to augment the training set of the summarization models, the performance on the noisy test set improves by up to $16\%$, indicating enhanced robustness against ASR errors. In summary, our contributions leading to the robustness are:
@@ 2 Related Work
Summarizing human conversations or speech has been a longstanding challenge that has received significant attention over the years (Hori and Furui 2003; Liu and Hakkani-Tür 2011). An important subset of these endeavors involves ASR-based summarization, where automatic speech recognition systems are used to transcribe spoken dialogues before summarization (Zhong et al. 2022). Recently, this challenge has extended to clinical applications, including the summarization of patient-doctor conversations (Krishna et al. 2021), utilizing LLMs as tools for text comprehension (Le-Duc et al. 2024).
((a))
((b))
@@
Data augmentation (DA) is crucial for enhancing model performance by diversifying training examples. Traditional DA methods in NLP, such as paraphrasing (Sharma et al. 2023), back-translation (Sugiyama and Yoshinaga 2019), and noise injection (Wang et al. 2018), have been effective in some scenarios, but may not fully capture the complexity of real-world variations. Recent LLM advancements have revolutionized DA by using models like GPT-3 and GPT-4 to generate high-quality synthetic data. For instance, Chintagunta et al. (2021) use in-context learning with LLMs to create synthetic medical dialogue summaries, while Dialogic (Li et al. 2022) employs LLMs to generate annotated dialogues with automatic verification and revision. DA is also used to improve robustness against adversarial examples, as used byLiu et al. (2020), who apply adversarial training to make sentiment analysis models more resilient to noise and variations. Our method differs by addressing the challenges of augmenting medical dialogue data where real ASR transcripts are scarce. Instead of relying on heuristic corruptions that don’t mimic ASR errors accurately, we use LLMs to generate synthetic dialogues that realistically replicate ASR errors, ensuring the augmented training data better aligns with real-world test conditions.
@@ 3.1 Error Generation using In-context Learning
To generate synthetic noisy dialogues, we first use a specialised system prompt to instruct the LLMs about the error generation task. Later we form the in-context examples by pairing sentences from human-transcribed clean dialogues with their ASR-generated counterparts. These examples are subsequently provided to the LLMs to convey the characteristics of ASR errors that we are trying to mimic. An example query that illustrates our structure when prompting the noise-generating LLMs is displayed in the following colored box. As seen from the example, the clean sentences are given as inputs and the responses are expected to contain erroneous versions as if they were obtained through ASR models. The curly braces surrounding certain substrings are a part of our error tagging system, which is detailed later in Section 3.2, and they denote which words or substrings are transcribed wrongly by the ASR models and how. For instance, in the example provided, the word “Tylenol” is wrongly transcribed as “tie-and-all”, highlighting the common issue of ASR systems confusing tokens with similar pronunciation. Lastly, the input at the end contains the actual clean sentence that we aim to corrupt. The words to be corrupted are also indicated to the LLM through the use of curly braces.
@@ 4 Experiment Settings
We utilized the Whisper tiny, Whisper large (Radford et al. 2023), and Wav2vec2-base (Baevski et al. 2020) ASR models to generate transcriptions of medical dialogues.
@@ 5 Preliminary Experiments
In this section, we present preliminary experiments that lay the foundation for our work. First, we conduct a motivating study that verifies how errors introduced by ASR systems negatively impact the performance of a downstream medical dialogue summarization task. Subsequently, we perform an analysis that suggests different ASR models exhibit distinct error profiles.
5.1 ASR noise harms medical report quality
Our findings, as displayed in Table 1, reveal that ASR noise can significantly degrade the quality of the generated summaries, especially when using small ASR models. Specifically, using Wav2vec2-base generated transcripts instead of human annotated ones lead to a noticeable reduction in the domain-specific entity F1 score of $23\%$ ($22.11\rightarrow 16.99$). While using larger models like Whisper-large exhibits comparable downstream task performance as using ground truth dialogues, deployment area of such large-scale models is limited due to the computational resource requirements they demand. Such a difference in the DSE overlap measure indicates a loss in accurately capturing critical medical entities. The ROUGE-L scores also decline, reflecting reduced textual overlap and coherence between the generated and reference summaries. Moreover, the BERT Scores drop, suggesting a decrease in the semantic similarity to the reference summaries. These results underscore the challenges of downstream tasks, such as dialogue summarisation, face, arising from erroneous ASR transcripts.
@@ 5.1 ASR noise harms medical report quality
We also explored applying few-shot denoising on ASR transcribed dialogues, which is denoted as “+ Denoising”. However, the results show that this method does not recover the summarization performance, highlighting the limitations of traditional post-processing based noise reduction techniques in this context.
@@ 6.2 LLM-produced synthetic errors are realistic
Llama-3-8B Summarization Mistral-7B Summarization
Source Transcripts ASR Random LLM-generated (MEDSAGE) ASR Random LLM-generated (MEDSAGE)
Llama-3-8B Mistral-7B Llama-3-8B Mistral-7B
F1
Wav2vec2-base 16.99 11.24 17.35 16.54 17.06 11.73 15.95 15.05
Whisper-tiny 19.86 12.83 18.02 17.91 18.04 15.38 18.63 18.10
Whisper-large 21.42 18.84 20.35 22.02 24.06 22.28 23.82 21.40
RougeL
Wav2vec2-base 11.52 8.50 11.47 11.18 11.09 8.32 11.70 10.34
Whisper-tiny 11.65 9.88 11.92 10.91 11.91 9.93 12.60 10.51
Whisper-large 13.67 11.98 12.52 12.70 13.27 11.82 13.47 11.39
Bert
Wav2vec2-base 52.46 49.59 52.87 52.71 50.40 44.61 51.78 49.56
Whisper-tiny 52.73 50.80 53.48 51.57 51.97 48.17 52.47 48.67
Whisper-large 54.23 52.07 53.70 53.59 53.29 50.09 53.64 49.94
Table 3: Summarization qualities for ASR and LLM-generated dialogue transcripts. Random stands for augmenting clean dialogues using random insertion, deletion and substitutions. The numbers in the ASR columns are colored blue to indicate reference values. The summarization qualities from synthetic dialogues closest to real ASR values are shown in bold.
Qualitative Comparison
To illustrate that synthetic dialogues produced by our MEDSAGE method properly mimics the errors caracteristics that real ASR transcriptions exhibit, we present randomly selected snippets from doctor-patient conversations in Figure 4.
@@
Figure 4: Comparison of an ASR transcription and its LLM-generated counterpart produced by MEDSAGE. The words highlighted in yellow are errors at common word indexes among both transcripts, while those that are highlighted in red and green indicate unique errors of ASR and MEDSAGE, respectively.
The commonalities between the nature of errors present in both utterances, such as phonetic confusions underscore the accuracy of the noise injections. For instance, the words ”white spots” are confused with ”whish spits”. This qualitative evidence supports our quantitative findings, demonstrating that LLM-generated synthetic noise can replicate the nuanced flaws typically seen in ASR outputs.
@@
To further substantiate the realism of LLM-generated synthetic noise, we conduct quantitative comparisons between the errors in synthetic dialogues and those in ASR-transcribed dialogues using the aforementioned metrics. First, we compute pairwise similarities, w.r.t. three evaluation metrics, among ASR transcripts and synthetic dialogues. As shown in the similarity matrices in Figure 3, the diagonal entries contain the highest similarity scores, indicating that the greatest overlap is between ASR transcripts and the corresponding synthetic dialogues generated to mimic them. This pattern suggests that the LLM-generated synthetic dialogues are successful in accurately replicating the specific error characteristics of the ASR transcripts they were designed to imitate.
Additionally, we provide an indirect comparison based on the summarization qualities each dialogue transcript yields. In this experiment, we also include dialogues corrupted with random deletions as well as insertions and substitutions from the NLTK words corpus (Bird, Klein, and Loper 2009) and name this baseline as Random. As shown in Table 3, the summarization qualities of LLM-generated synthetic dialogues closely align with those of the ASR transcripts, which are indicated in blue as reference values. This further supports the effectiveness of our method in producing realistic synthetic data that mimics ASR errors.
@@
6.3 Error tagging system enables controllable generation
Lastly, we validate that our method is capable of adjusting the injected noise rate through the use of corruption tags, as detailed in the methodology section. To this end, we insert corruption tags at various rates to simulate increasing amounts of noise. We then record the resulting transcription quality and summarization performance for each noise level.
As depicted in Figure 5, we observe a clear trend: as the noise rate increases, the quality of the generated summaries decreases. This is evidenced by a gradual drop in the summarization quality quantified by the scores on various metrics. This analysis demonstrates the controllability of our noise generation method, which is essential for adapting the synthetic noise generation to errors made by various different ASR models with ease.
@@ 7 Conclusion
Appendix A: Computing Infrastructure
The specifications of the system used to run computational experiments are:
  • Ubuntu 22.04.4 LTS
  • NVIDIA A100 80GB GPUs
  • Intel(R) Xeon(R) Silver 4314 CPU @ 2.40GHz
Appendix B: Additional Experiment Details & Results
Here we present additional qualitative and quantitative experimental results that extend beyond those given in the main paper.
Huggingface Models Used in Experiments
For the experiments resulting in Table 2 of the main paper, we fine-tuned the following model checkpoints retrieved from Hugging Face:
  • meta-llama/Meta-Llama-3-8B-Instruct
  • mistralai/Mistral-7B-Instruct-v0.2
  • unsloth/gemma-7b-it
For the remaining experiments, we used AWQ quantized checkpoints of the same models:
  • solidrust/Meta-Llama-3-8B-Instruct-hf-AWQ
  • TheBloke/Mistral-7B-Instruct-v0.2-AWQ
For inference, we primarily used the Text Generation Inference (TGI) framework [^1] https://github.com/huggingface/text-generation-inference. For fine-tuning, we utilized Unsloth [^2] https://github.com/Unsloth/unsloth.
Extended Evaluation Using Other Metrics
The experimental evaluation in the main paper primarily reports F1 score over domain-specific entities, Rouge-L, and Bert score. Here, we extend this evaluation by presenting results using additional metrics such as Precision, Recall, Rouge-1, and Rouge-2.
Figure 6: Precision, Recall, and F1 score across different noise levels. The shaded regions represent the standard deviation (STD) across multiple runs.
Figure 7: Rouge-1, Rouge-2, Rouge-L, and Rouge-Lsum scores across different noise levels. The shaded regions represent the standard deviation (STD) across multiple runs.
Figures 6 and 7, also included in the attachment, illustrate how these additional metrics behave under varying noise levels. It is observed that these metrics, such as Precision, Recall, Rouge-1, and Rouge-2, closely correlate with the Rouge-L and F1 scores reported in the main paper. Specifically, as noise increases, a noticeable decline is observed across all metrics.
Snippet 1:
(Whisper-tiny ASR + Denoising)
Patient: Comes and goes, yeah.
Doctor: Is there a pain, somewhere else, moving down between your shoulders?
Patient: Um, no, actually… just possibly my stomach.
(Whisper-tiny ASR)
Patient: comes and goes.
Doctor: There’s a pain move anywhere else, with on-between your back.
Patient: Um, no, just maybe my stomach.
(Ground truth)
Patient: Uh, it comes and goes.
Doctor: Come and go. Does the pain move anywhere else, for example towards your back?
Patient: Uh…no, just maybe my stomach.
Snippet 2:
(Whisper-tiny ASR + Denoising)
Doctor: OK, an’…an’ yep, you did mention having some pain in yer tummies too, didn’t ya? Whereabouts is the pain?
Patient: Ah, yes… actually, I don’t think this has got anything to do with my elbow… I mean, aren’t we talking about something else? Like, ah, in my lower abdomen… specifically, towards one side…?
Doctor: One side…um, what did you mean by ”which” side?
(Whisper-tiny ASR)
Doctor: Okay, and you mentioned you’ve had some pain in your tummy as well. Yep, where about as is the pain?
Patient: Yeah, we’re exactly in my lower abdomen, so like, yeah, just to one side.
Doctor: one side of what side is that?
(Ground truth)
Doctor: Okay. And you mentioned you’ve had some pain in your tummy as well. Whereabouts is the pain, exactly?
Patient: Yep. So in my lower abdomen, so, uh, like, um…yeah, just to one side.
Doctor: One side. And what side is that?
Figure 8: Comparison of dialogue transcripts showing hallucinated words by the denoising model, highlighted in red. Yellow highlights indicate the correct words in the ASR and ground truth transcriptions corresponding to the hallucinated substrings in the denoising output.
Denoising Models can Hallucinate
As outlined in the main paper, one significant challenge associated with using LLMs for ASR error correction, without fine-tuning, is their inclination to hallucinate, or generate content that is not present in the original input. This issue can severely degrade the quality of medical reports that are generated based on dialogues that are pre-processed by denoising models. Figure 8 provides qualitative examples that highlight this problem. These examples illustrate how a denoising model, designed to clean up noisy ASR transcriptions, introduces additional errors in the form of hallucinations.
In Snippet 1, for example, the denoising model generates the phrase ”your shoulders,” which is not present in either the ASR output or the ground truth transcript. Similarly, in Snippet 2, the denoising model introduces a segment that is non-existent in the actual converstaion: ”actually, I don’t think this has got anything to do with my elbow… I mean, aren’t we talking about something else?” These hallucinations can lead to inaccurate representations of the dialogue, which is particularly problematic when the text is later used for summarization or decision-making in clinical settings.
Diversity of Error Styles Across ASR Models
The analysis of ASR errors reveals that these errors are often rooted in phonetic confusions, where the model misinterprets similar-sounding words or phrases. However, the specific types of phonetic confusions are unique to the ASR model employed, as demonstrated in the examples provided in Figure 9. For instance, the Whisper Tiny model shows errors such as ”diure” instead of ”diarrhea,” while the Wav2vec2 model struggles with more severe distortions like ”TWOLLE” instead of ”toilet.” This variability of error styles across ASR models underscores the challenge of employing a one-fits-all synthetic dialogue generation approach. To address this, our LLM-based synthetic data generation framework incorporates in-context examples that capture the distinctive error patterns of each ASR model. By providing the LLM with these targeted examples, we ensure that it qualitatively understands the specific types of ASR errors we aim to mimic.
Whisper Tiny (ASR) Doctor: Sorry to hear that. And when you say diure, what do you mean by diure? Do you mean you’re going to tell a more often or are your schools more loose? Yes. Patient: Yeah, so it’s like loose and watery stall going to the toilet quite often and like some pain in my like lower stomach. Okay Doctor: And how many times a day are you going? Let’s see, a lot of a couple of days.
Wav2vec2 (ASR) Doctor: SO YO HEAR BATS AND WE YOU SAID DIRE AR WHAT DO YOU MEAN BY DIRECT YOU MEAN YOU GAINST THE TWOLLE MORE OFTEN OR ARE YOUR STOOLS MORE LOOSE YE Patient: YEA SO IT’S LIKE LOOSE AND MORDERY STOL GOING TO THE TOILET QUITE OFTEN LA AND LIKE SOME PAIN IN MY LIKE LOWER STOMACH LIK Doctor: LA N E AND HAWMY TIMES A DAY ARE YOU GOING LETS SEE ER LOST COUPLE DAYS
Whisper Large (ASR) Doctor: Sorry to hear that. And when you say diarrhea, what do you mean by diarrhea? Do you mean you go into the toilet more often or are your stools more loose? Patient: Yeah, so it’s like loose and watery stool going to the toilet quite often and like some pain in my like lower stomach. Okay Doctor: How many times a day are you going?
Ground Truth Doctor: Sorry to hear that. Um, and and when you say diarrhea, what’d you mean by diarrhea? Do you mean you’re going to the toilet more often? Or are your stools more loose? Patient: Yeah, so it’s like loose and watery stool, going to the toilet quite often, uh and like some pain in my, like, lower stomach? Doctor: Okay. And how many times a day are you going, let’s say, in the last couple of days?
Figure 9: Comparison of dialogue transcripts generated by different ASR models and the ground truth. The figure illustrates how phonetic confusions are common across ASR models, while each model exhibits a unique error style.

### v3

@@ 1 Introduction
Automatic Speech Recognition (ASR) (Yu and Deng 2016) is the task of transcribing speech signals into text, enabling a wide range of applications from voice-activated assistants to automated customer service systems. ASR systems significantly aid various downstream tasks such as dialogue summarization (Liu et al. 2019; Zhong et al. 2022), where the goal is to distill key information from spoken interactions. However, errors introduced by ASR systems can degrade the performance of these summarization tasks (Li et al. 2014; Guo et al. 2024), which limits their application in high-stake domains where correctness of the summaries is important. The synthesis of Electronic Medical Records (EMRs) from doctor-patient dialogues (Krishna et al. 2021) is among such summarization tasks where accuracy is critical. Here, ASR errors can lead to inaccuracies in transcribing clinical terms, medication, or procedure names, resulting in erroneous medical notes (Hodgson and Coiera 2016), which can lead to misdiagnosis, incorrect treatment plans, and potentially harmful patient outcomes.
One way to improve dialogue summarization is to improve the ASR systems. However, this requires large amounts of supervised data (Radford et al. 2022), which is often unavailable in the healthcare domain due to privacy and ethical concerns. Consequently, clinical summarization systems often treat ASR systems as black boxes, mandating that overall improvements must arise either from post-processing ASR outputs, or making down-stream methods robust to ASR errors. Recent works propose post-processing ASR outputs using LLMs to correct erroneous ASR transcripts (Radhakrishnan et al. 2023; Bai et al. 2024). Nonetheless, based on prior studies, using prompting techniques to clear noise only proves effective only when using LLMs of large sizes, typically those that exceed 100B parameters (Yang et al. 2023). Moreover, LLMs are prone to hallucinations (Maynez et al. 2020; Nagar et al. 2024), which can introduce irrelevant symptoms or medication names into the dialogue transcript, potentially degrading summarization quality while attempting to clear noise.
@@
Our experimental evaluation reveals that large language models (LLMs) are effective noise modelers, capable of producing similar errors to those produced by ASR. This is evident from the qunatitative and qualitative similarities between the word error profiles of the synthetic dialogues we generate and actual ASR dialogue transcripts. Such similarity also reflects on the downstream summarization performance, mirroring the performance drop caused by ASR errors. Lastly, when we utilize these synthetic noisy dialogues to augment the training set of the summarization models, the performance on the noisy test set improves by up to $16\%$, indicating enhanced robustness against ASR errors. In summary, our contributions leading to the robustness are:
@@ 2 Related Work
Summarizing human conversations or speech has been a longstanding challenge that has received significant attention over the years (Hori and Furui 2003; Liu and Hakkani-Tür 2011). An important subset of these endeavors involves ASR-based summarization, where automatic speech recognition systems are used to transcribe spoken dialogues before summarization (Zhong et al. 2022). Recently, this challenge has extended to clinical applications, including the summarization of patient-doctor conversations (Krishna et al. 2021), utilizing few-shot (Le-Duc et al. 2024) or fine-tuned (Schlegel et al. 2023) LLMs as tools for text comprehension.
@@
Data augmentation (DA) is crucial for enhancing model performance by diversifying training examples. Traditional DA methods in NLP, such as paraphrasing (Sharma et al. 2023; Li et al. 2023), back-translation (Sugiyama and Yoshinaga 2019), and noise injection (Wang et al. 2018), have been effective in some scenarios, but may not fully capture the complexity of real-world variations. Recent LLM advancements have revolutionized DA by using models like GPT-3 and GPT-4 to generate high-quality synthetic data. For instance, Chintagunta et al. (2021) use in-context learning with LLMs to create synthetic medical dialogue summaries, while Dialogic (Li et al. 2022) employs LLMs to generate annotated dialogues with automatic verification and revision. Similarly, Liu et al. (2024) leverage topic-focused summarization and domain adaptation with LLMs to generate personalized medical dialogues. DA is also used to improve robustness against adversarial examples, as used by Liu et al. (2020), who apply adversarial training to make sentiment analysis models more resilient to noise and variations. Our method differs by addressing the challenges of augmenting medical dialogue data where real ASR transcripts are scarce. Instead of relying on heuristic corruptions that don’t mimic ASR errors accurately, we use LLMs to generate synthetic dialogues that realistically replicate ASR errors, ensuring the augmented training data better aligns with real-world test conditions.
@@ 3.1 Error Generation using In-context Learning
To generate synthetic noisy dialogues, we first use a specialised system prompt to instruct the LLMs about the error generation task. Later we form the in-context examples by pairing sentences from human-transcribed clean dialogues with their ASR-generated counterparts. These examples are subsequently provided to the LLMs to convey the characteristics of ASR errors that we are trying to mimic. An example query that illustrates our structure when prompting the noise-generating LLMs is displayed in the following box.
@@
As seen from the example, the clean sentences are given as inputs and the responses are expected to contain erroneous versions as if they were obtained through ASR models. The curly braces surrounding certain substrings are a part of our error tagging system, which is detailed later in Section 3.2, and they denote which words or substrings are transcribed wrongly by the ASR models and how. For instance, in the example provided, the word “Tylenol” is wrongly transcribed as “tie-and-all”, highlighting the common issue of ASR systems confusing tokens with similar pronunciation. Lastly, the input at the end contains the actual clean sentence that we aim to corrupt. The words to be corrupted are also indicated to the LLM through the use of curly braces.
@@ 4 Experiment Settings
We used the Whisper tiny/ large(Radford et al. 2023), and Wav2vec2-base (Baevski et al. 2020) ASR models to generate transcriptions of medical dialogues.
@@ 5 Preliminary Experiments
In this section, we present preliminary experiments that lay the foundation for our work. First, we conduct a motivating study that verifies how ASR errors impair the performance of a downstream medical dialogue summarization task. Subsequently, we perform an analysis that suggests different ASR models exhibit distinct error profiles.
@@
5.1 ASR noise harms medical report quality
Our findings, as displayed in Table 1, reveal that ASR noise can significantly degrade the quality of the generated summaries, especially when using small ASR models. Specifically, using Wav2vec2-base generated transcripts instead of human annotated ones lead to a noticeable reduction in the domain-specific entity F1 score of $23\%$ ($22.11\rightarrow 16.99$). While using larger models like Whisper-large exhibits comparable downstream task performance as using ground truth dialogues, deployment area of such large-scale models is limited due to the computational resource requirements they demand. Such a difference in the DSE overlap measure indicates a loss in accurately capturing critical medical entities. The ROUGE-L scores also decline, reflecting reduced textual overlap and coherence between the generated and reference summaries. Moreover, the BERT Scores drop, suggesting a decrease in the semantic similarity to the reference summaries. These results underscore the challenges of downstream tasks, such as dialogue summarisation, face, arising from erroneous ASR transcripts. We also explored applying few-shot denoising on ASR transcribed dialogues, which is denoted as “+ Denoising”. However, the results show that this method does not recover the summarization performance, highlighting the limitations of traditional post-processing based noise reduction techniques in this context.
@@ 6.2 LLM-produced synthetic errors are realistic
Figure 3: Comparison of an ASR transcription and its LLM-generated counterpart produced by MEDSAGE. The words highlighted in yellow are errors at common word indexes among both transcripts, while those that are highlighted in red and green indicate unique errors of ASR and MEDSAGE.
Qualitative Comparison
To illustrate that synthetic dialogues produced by our MEDSAGE method properly mimics the errors characteristics that real ASR transcriptions exhibit, we present randomly selected snippets from doctor-patient conversations in Figure 3. The commonalities between the nature of errors present in both utterances, such as phonetic confusions underscore the accuracy of the noise injections. For instance, the words ”white spots” are confused with ”whish spits”. This qualitative evidence supports our quantitative findings, demonstrating that LLM-generated synthetic noise can replicate the nuanced flaws typically seen in ASR outputs.
Llama-3-8B Summarization
Source Transcripts ASR Random LLM-generated (MEDSAGE)
Llama-3-8B Mistral-7B
F1
Wav2vec2-base 16.99 11.24 17.35 16.54
Whisper-tiny 19.86 12.83 18.02 17.91
Whisper-large 21.42 18.84 20.35 22.02
RougeL
Wav2vec2-base 11.52 8.50 11.47 11.18
Whisper-tiny 11.65 9.88 11.92 10.91
Whisper-large 13.67 11.98 12.52 12.70
Bert
Wav2vec2-base 52.46 49.59 52.87 52.71
Whisper-tiny 52.73 50.80 53.48 51.57
Whisper-large 54.23 52.07 53.70 53.59
Mistral-7B Summarization
Source Transcripts ASR Random LLM-generated (MEDSAGE)
Llama-3-8B Mistral-7B
F1
Wav2vec2-base 17.06 11.73 15.95 15.05
Whisper-tiny 18.04 15.38 18.63 18.10
Whisper-large 24.06 22.28 23.82 21.40
RougeL
Wav2vec2-base 11.09 8.32 11.70 10.34
Whisper-tiny 11.91 9.93 12.60 10.51
Whisper-large 13.27 11.82 13.47 11.39
Bert
Wav2vec2-base 50.40 44.61 51.78 49.56
Whisper-tiny 51.97 48.17 52.47 48.67
Whisper-large 53.29 50.09 53.64 49.94
Table 3: Summarization qualities for ASR and LLM-generated dialogue transcripts. Random stands for augmenting clean dialogues using random insertion, deletion, and substitutions. The ASR columns are colored blue to indicate reference values. The summarization qualities from synthetic dialogues closest to real ASR are shown in bold.
@@
To further substantiate the realism of LLM-generated synthetic noise, we conduct quantitative comparisons between the errors in synthetic dialogues and those in ASR-transcribed dialogues using the aforementioned metrics. First, we compute pairwise similarities, w.r.t. three evaluation metrics, among ASR transcripts and synthetic dialogues. As shown in the similarity matrices in Figure 4, the diagonal entries contain the highest similarity scores, indicating that the greatest overlap is between ASR transcripts and the corresponding synthetic dialogues generated to mimic them. This pattern suggests that the LLM-generated synthetic dialogues are successful in accurately replicating the specific error characteristics of the ASR transcripts they were designed to imitate. Additionally, we provide an indirect comparison based on the summarization qualities each dialogue transcript yields. In this experiment, we also include dialogues corrupted with random deletions as well as insertions and substitutions from the NLTK words corpus (Bird, Klein, and Loper 2009) and name this baseline as Random. As shown in Table 3, the summarization qualities of LLM-generated synthetic dialogues closely align with those of the ASR transcripts, which are indicated in blue as reference values. This further supports the effectiveness of our method in producing realistic synthetic data.
6.3 Error tagging system enables controllable generation
Lastly, we validate that our method is capable of adjusting the injected noise rate through the use of corruption tags, as detailed in the methodology section. To this end, we insert corruption tags at various rates to simulate increasing amounts of noise. We then record the resulting transcription quality and summarization performance for each noise level. As depicted in Figure 5, we observe a clear trend: as the noise rate increases, the quality of the generated summaries decreases. This is evidenced by a gradual drop in the summarization quality quantified by the metric scores. This demonstrates the controllability of our noise generation method, which is essential for adapting the synthetic noise generation to errors made by various different ASR models.
@@ 7 Conclusion
Acknowledgements
Viktor is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.